OpenStudy (anonymous):

10k^2 divided by 3 times h^5 divided by 20k^5

OpenStudy (anonymous):

\[10k^2\div3\times h^5\div 20k^5\]

OpenStudy (diyadiya):

\[\large \frac{10k^2}{3} \times \frac{h^5}{20k^5}\]

OpenStudy (anonymous):

r u there?

OpenStudy (diyadiya):

\[\large \frac{10k^2 \times h^5}{3 \times 20k^5} \]\[\frac{10}{20}= \frac{1}{2}\]\[\Large \frac{k^2}{k^5}= k^{2-5}=k^{-3}\]

OpenStudy (diyadiya):

Remember \[\Large \frac{a^m}{a^n}= a^{m-n}\]

OpenStudy (diyadiya):

Also remember \[\Large a^{-m}= \frac{1}{a^m}\]

OpenStudy (diyadiya):

\[\Large \frac{ h^5}{6k^3}\]

OpenStudy (diyadiya):

Did you understand ?

OpenStudy (anonymous):

i dont understand! what is the answer?

OpenStudy (diyadiya):

Did you undertand this ?